OpenStudy (anonymous):

a projectile is fired at a speed of 20.0 m/s upward at an angle of 53.0 relative to the horizontal x-axis. ignoring air resistance, calculate the time the projectile is in air?

OpenStudy (anonymous):

I think this can be done like this, since time of flight when projectile fired from ground with horizontal = \[(2u \sin \theta)\div g\] (when air resistance is neglected). where u= speed of projectile \[\theta =\]angle g= acceleration due to gravity which is 9.8 m/s^2 now put these values in formula and solve.